# Contest for system observability as an infinitely repeated game

### v2

@@ 3.2 The best response set of the defender
which is equal to $\arg\max_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Im}\Omega^{T}$,where $\Omega^{T}\!=[H^{\top}\!,\!\left(N+B_{a}^{\prime}U_{a}\right)^{\top}\!H^{\top},\!\cdots,\left(N+B_{a}^{\prime}U_{a}\right)^{(n-1)\top}\!H^{\top}]$ can be viewed as controllability matrix of the dual system of (9):
$\dot{\bar{z}}=\bar{A}\bar{z}+\bar{B}\bar{u},$ (17)
where $\bar{A}=(N+B_{a}^{\prime}U_{a})^{\top},\bar{B}=H^{\top}$. Thus the best response set (16) becomes
$\bar{B}^{*}=\arg\max_{\bar{B}}\operatorname{dim\,Im}\left[\begin{array}[]{lll}\bar{B},\!\!&\bar{A}\bar{B},\cdots,\bar{A}^{n-1}\bar{B}\end{array}\!\!\right].$ (18)
The problem becomes how to choose $\bar{B}$ to make the dual system controllable. Find similar transformation matrix $T\in\mathbb{R}^{n\times n}$ which makes $\bar{A}$ become Jordan normal form $J$, i.e., $J=T^{-1}\bar{A}T$. And $\bar{B}$ becomes $\hat{B}=T^{-1}\bar{B}$. For the above $J$, let its $l$ eigenvalues be: $\lambda_{1}$ (algebraic multiplicity: $\sigma_{1}$, geometric multiplicity: $\alpha_{1}$), $\lambda_{2}$ (algebraic multiplicity: $\sigma_{2}$, geometric multiplicity: $\alpha_{2}$), $\cdots$, $\lambda_{l}$ (algebraic multiplicity: $\sigma_{l}$, geometric multiplicity: $\alpha_{l}$). Assume $\lambda_{i}\neq\lambda_{j},\forall i\neq j$ and $\sigma_{1}+\sigma_{2}+\cdots+\sigma_{l}=n$.
Thus we have
$J=J\left(\lambda_{1}\right)\oplus J\left(\lambda_{2}\right)\oplus\cdots\oplus J\left(\lambda_{l}\right),$ (19)
$\hat{B}=\left[\hat{B}_{1}^{\top},\hat{B}_{2}^{\top},\cdots,\hat{B}_{l}^{\top}\right]^{\top},$ (20)
@@
$J(\lambda_{i})=\begin{bmatrix}{J}_{1}(\lambda_{i})&&&\\ &{J}_{2}(\lambda_{i})&&\\ &&\ddots&\\ &&&{J}_{\alpha_{i}}(\lambda_{i})\end{bmatrix},\hat{B}_{i}=\begin{bmatrix}\hat{{B}}_{i1}\\ \hat{{B}}_{i2}\\ \vdots\\ \hat{{B}}_{i\alpha_{i}}\end{bmatrix},$
@@
$J_{k}(\lambda_{i})=\begin{bmatrix}\lambda_{i}&1&&&\\ &\lambda_{i}&1&&\\ &&\ddots&\ddots&\\ &&&\ddots&1\\ &&&&\lambda_{i}\end{bmatrix},\hat{{B}}_{ik}=\begin{bmatrix}\hat{{b}}^{ik}_{1}\\ \hat{{b}}^{ik}_{2}\\ \vdots\\ \hat{{b}}^{ik}_{r_{ik}-1}\\ \hat{{b}}^{ik}_{r_{ik}}\end{bmatrix},$
for $k=1,2,\cdots,\alpha_{i}$, where $J_{k}(\lambda_{i})\in\mathbb{R}^{r_{ik}\times r_{ik}},\hat{{B}}_{ik}\in\mathbb{R}^{r_{ik}\times m}$. $\sum_{k=1}^{\alpha_{i}}r_{ik}=\sigma_{i}$. Recall the following Lemma.
Lemma 2.
@@
$\operatorname{rank}\left[\hat{b}_{r_{i1}}^{i1\top},\hat{b}_{r_{i2}}^{i2\top},\cdots,\hat{b}_{r_{i\alpha_{i}}}^{i\alpha_{i}\top}\right]=\alpha_{i},\forall i=1,2,\cdots,l,$ (21)
which means the last rows of $\hat{{B}}_{i1},\hat{{B}}_{i2},\cdots,\hat{{B}}_{i\alpha_{i}}$ are linearly independent.
However, if there is a geometric multiplicity of an eigenvalue that exceeds the number of columns of $\hat{B}$, i.e., $\exists\alpha_{i}>m$, it is impossible for the system to be fully controllable. In this case, we give a condition to maximize the controllable subspace.
@@
$\hat{B}\in\mathbb{R}^{n\times m}$ maximizes the dimension of the controllable subspace if and only if the last rows of $\hat{{B}}_{i1},\hat{{B}}_{i2},\cdots,\hat{{B}}_{i\alpha_{i}}$ satisfy
$\left\{\begin{array}[]{cc}\operatorname{rank}\left[\hat{b}_{r_{i1}}^{i1\top},\hat{b}_{r_{i2}}^{i2\top},\cdots,\hat{b}_{r_{i\alpha_{i}}}^{i\alpha_{i}\top}\right]=\alpha_{i},\,\text{for}\,\{i\mid m\geq\alpha_{i}\},\\ \\ \operatorname{rank}\left[\hat{b}_{r_{ij_{1}}}^{ij_{1}\top},\hat{b}_{r_{ij_{2}}}^{ij_{2}\top},\cdots,\hat{b}_{r_{ij_{m}}}^{ij_{m}\top}\right]=m,\,\text{for}\,\{i\mid m<\alpha_{i}\},\end{array}\right.$ (22)
where $\{j_{1},j_{2},\cdots,j_{m}\}\subset\{1,2,\cdots,\alpha_{i}\}$ are the subscripts of $m$-th largest Jordan blocks for eigenvalue $\lambda_{i}$.
Proof 3.2.
For $m\geq\alpha_{i}$, (sufficiency and necessity) according to Lemma 2, it is the sufficient and necessary condition for complete controllability. Thus the dimension of controllable subspace is maximized.
For $m<\alpha_{i}$, (sufficiency) if $\operatorname{rank}\left[\hat{b}_{r_{ij_{1}}}^{ij_{1}\top},\hat{b}_{r_{ij_{2}}}^{ij_{2}\top},\cdots,\hat{b}_{r_{ij_{m}}}^{ij_{m}\top}\right]$ $=m$, the dimension of controllable subspace $\operatorname{dim\,Im}\Omega^{\top}=\sum_{p=1}^{m}r_{ij_{p}}.$ Because $\{j_{1},j_{2},\cdots,j_{m}\}$ are the subscripts of $m$-th largest Jordan blocks for eigenvalue $\lambda_{i}$, the dimension of controllable subspace is maximized. (Necessity) Given the symmetry in the argument, the necessity part is omitted.
With $\hat{B}$ which satisfies Proposition 3.1, the best response for the defender is
@@
where $T$ is the transformation matrix making $(N+B_{a}^{\prime}U_{a})^{\top}$ become Jordan normal form. Based on Proposition 3.1, we can obtain the lower bound of the value function minimized by the defender.
Corollary 3.3.
For ${H}\in\mathbb{R}^{m\times n}$, define the set $\mathcal{I}$ which includes subscripts for the eigenvalues of $(N+B_{a}^{\prime}U_{a})$ whose geometric multiplicity is larger than $m$, i.e., =$\mathcal{I}=\{i\mid\alpha_{i}>m\}$. Then $\forall U_{a}\in\mathbb{R}^{k\times n}$,
$&\underset{{H}\in\mathbb{R}^{m\times n}}{\min}\!\operatorname{dim\,Ker}\Omega(U_{a},H)\!=\!n\!-\!\underset{{q}\in\mathcal{I}}{\min}(\sum_{p=1}^{m}r_{qj_{p}}),\,\text{for }\mathcal{I}\neq\emptyset,\\ &\underset{{H}\in\mathbb{R}^{m\times n}}{\min}\operatorname{dim\,Ker}\Omega(U_{a},H)=0,\,\text{for }\mathcal{I}=\emptyset,\end{aligned}\right.$ (24)
where $\{r_{qj_{1}},r_{qj_{2}},\cdots,r_{qj_{m}}\}$ are the dimensions of $m$-th largest Jordan blocks for eigenvalue $\lambda_{q}$.
Based on Proposition 3.1, we can also give a specific way to construct $\hat{{B}}$.
Corollary 3.4.
If the last rows of $\hat{{B}}_{i1},\hat{{B}}_{i2},\cdots,\hat{{B}}_{i\alpha_{i}}$ satisfy
$\left[\!\begin{array}[]{c}\hat{b}_{r_{i1}}^{i1}\\ \hat{b}_{r_{i2}}^{i2}\\ \vdots\\ \hat{b}_{r_{i\alpha_{i}}}^{i\alpha_{i}}\end{array}\!\right]\!\!=\!\!\left\{\begin{array}[]{cc}\left[I_{\alpha_{i}}\ \textbf{0}_{\alpha_{i}\times\left(m-\alpha_{i}\right)}\right],\,\text{for}\,\{i\mid m\geq\alpha_{i}\},\\ \\ \left[\delta_{\alpha_{i}}^{j_{1}},\delta_{\alpha_{i}}^{j_{2}},\cdots,\delta_{\alpha_{i}}^{j_{m}}\right],\,\text{for}\,\{i\mid m<\alpha_{i}\},\end{array}\right.$ (25)
where  $\delta_{\alpha_{i}}^{j_{p}}=\operatorname{Col}_{j_{p}}\left(I_{\alpha_{i}}\right)$, $\{j_{1},j_{2},\cdots,j_{m}\}\subset\{1,2,\cdots,\alpha_{i}\}$ are the subscripts of $m$-th largest Jordan blocks for eigenvalue $\lambda_{i}$ and other rows of $\hat{{B}}_{i1},\hat{{B}}_{i2},\cdots,\hat{{B}}_{i\alpha_{i}}$ are zero rows, then the controllable subspace is maximized.
The calculation steps of $H^{*}$ based on Corollary 3.4 is summarized in Algorithm 2.
Algorithm 2 Algorithm for the defender to minimize the unobservable subspace
  Set $\bar{A}=(N+B_{a}^{\prime}U_{a})^{\top}.$ Calculate Jordan normal form of $\bar{A}:T^{-1}\bar{A}T=J$. Calculate geometric multiplicity $\alpha_{i}$ of $J$’s eigenvalue $\lambda_{i},i=1,2,\cdots,l$ ; Calculate dimensions $\{r_{i1},r_{i2},\cdots,r_{i\alpha_{i}}\}$ of Jordan blocks of $\lambda_{i},i=1,2,\cdots,l$; Calculate $\{j_{1},j_{2},\cdots,j_{m}\}$ which are the subscripts of $m$-th largest Jordan blocks for eigenvalue $\lambda_{i}$. Set $\hat{B}\in\mathbb{R}^{n\times m}$ to be an all-zeros matrix. Update $\hat{B}$ according to (25). Calculate $H=\hat{B}^{\top}T^{\top}$.
@@ 4 Equilibrium analysis
Based on the above best response sets, we further analyze the equilibrium of the game considering one-step optimality and two-step optimality respectively in the subsections 4.1 and 4.2. In the subsection 4.3, we give equilibrium analysis which is suitable for both cases.
@@ 4.1 Game based on one-step optimality
The infinitely repeated game considering one-step optimality means that players only consider the best of the current epoch. Define the best response sets for the attacker and the defender in this case as $BR1^{a}$ and $BR1^{d}$, i.e.,
$BR1^{a}=\arg\max_{{U_{a}}\in\mathbb{R}^{k\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H),$ (26)
$BR1^{d}=\arg\min_{{H}\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H).$ (27)
Fig. 2 shows the sequence of actions. The solutions of those two best response sets have been discussed in subsections 3.1 and 3.2, i.e.,
$BR1^{a}=\mathcal{F}(\mathcal{V}^{*}),$ (28)
where $\mathcal{F}(\mathcal{V}^{*})$ is the friend matrices of $\mathcal{V}^{*}$, and $\mathcal{V}^{*}$ is the maximal controlled invariant subspace in $\operatorname{Ker}H$. And
@@
where $T$ is the transformation matrix making $(N+B_{a}^{\prime}U_{a})^{\top}$ become Jordan normal form. Both sets are not single-valued maps and Algorithm 1 (or Algorithm 2) only determines a special $U_{a}$ (or $H$) in $BR1^{a}$ (or $BR1^{d}$). If a player chooses a different action in the best response set, the game result will be different.
Figure 2: Sequence of actions considering one-step optimality.
Define the Nash equilibrium.
Definition 2.
The strategy profile $(U_{a}^{*},H^{*})$ is said to be the Nash equilibrium of the one-step optimal game, if $U_{a}^{*}\in BR1^{a},H^{*}\in BR1^{d}$.
@@
Theorem 4.5.
The strategy profile $(U_{a}^{*},H^{*})$ is a Nash equilibrium of the one-step optimal game if and only if $\min_{{H}\in\mathbb{R}^{m\times n}}\!\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},\!H)\!=\!\operatorname{dim}\mathcal{V}^{*}(H^{*})\!=\!\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},H^{*})$.

### v3

@@ 3.2 The best response set of the defender
which is equal to $\arg\max_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Im}\Omega^{T}$,where
$\Omega^{T}\!=[H^{\top}\!,\!\left(N+B_{a}^{\prime}U_{a}\right)^{\top}\!H^{\top},\!\cdots,\left(N+B_{a}^{\prime}U_{a}\right)^{(n-1)\top}\!H^{\top}]$
can be viewed as the controllability matrix of the system (7)’s dual system: $\dot{\overline{z}}=\overline{A}\overline{z}+\overline{B}\overline{u}$, where $\overline{A}=(N+B_{a}^{\prime}U_{a})^{\top},\overline{B}=H^{\top}$. Thus the best response set (12) becomes $\overline{B}^{*}=\arg\max_{\overline{B}}\operatorname{dim\,Im}\left[\begin{array}[]{lll}\overline{B},\!\!&\overline{A}\overline{B},\cdots,\overline{A}^{n-1}\overline{B}\end{array}\!\!\right]$. The problem becomes how to choose $\overline{B}$ to make the dual system controllable. Find similar transformation matrix $T\in\mathbb{R}^{n\times n}$ which makes $\overline{A}$ become Jordan normal form $J$, i.e., $J=T^{-1}\overline{A}T$. And $\overline{B}$ becomes $\widehat{B}=T^{-1}\overline{B}$. For the above $J$, let its $l$ eigenvalues be: $\lambda_{1}$ (algebraic multiplicity: $\sigma_{1}$, geometric multiplicity: $\alpha_{1}$), $\lambda_{2}$ (algebraic multiplicity: $\sigma_{2}$, geometric multiplicity: $\alpha_{2}$), $\cdots,\lambda_{l}$ (algebraic multiplicity: $\sigma_{l}$, geometric multiplicity: $\alpha_{l}$). Assume $\lambda_{i}\neq\lambda_{j},\forall i\neq j$ and $\sigma_{1}+\sigma_{2}+\cdots+\sigma_{l}=n$. Thus we have
$J=J\left(\lambda_{1}\right)\oplus J\left(\lambda_{2}\right)\oplus\cdots\oplus J\left(\lambda_{l}\right),\,\widehat{B}=\left[\widehat{B}_{1}^{\top},\widehat{B}_{2}^{\top},\cdots,\widehat{B}_{l}^{\top}\right]^{\top},$
@@
$J(\lambda_{i})=\begin{bmatrix}{J}_{1}(\lambda_{i})&&&\\ &{J}_{2}(\lambda_{i})&&\\ &&\ddots&\\ &&&{J}_{\alpha_{i}}(\lambda_{i})\end{bmatrix},\,\widehat{B}_{i}=\begin{bmatrix}\widehat{{B}}_{i1}\\ \widehat{{B}}_{i2}\\ \vdots\\ \widehat{{B}}_{i\alpha_{i}}\end{bmatrix},$
@@
$J_{k}(\lambda_{i})=\begin{bmatrix}\lambda_{i}&1&&&\\ &\lambda_{i}&1&&\\ &&\ddots&\ddots&\\ &&&\ddots&1\\ &&&&\lambda_{i}\end{bmatrix},\,\widehat{{B}}_{ik}=\begin{bmatrix}\widehat{{b}}^{ik}_{1}\\ \widehat{{b}}^{ik}_{2}\\ \vdots\\ \widehat{{b}}^{ik}_{r_{ik}-1}\\ \widehat{{b}}^{ik}_{r_{ik}}\end{bmatrix},$
for $k=1,2,\cdots,\alpha_{i}$, where $J_{k}(\lambda_{i})\in\mathbb{R}^{r_{ik}\times r_{ik}},\widehat{{B}}_{ik}\in\mathbb{R}^{r_{ik}\times m}$, $\sum_{k=1}^{\alpha_{i}}r_{ik}=\sigma_{i}$. Recall the following Lemma which gives the controllability condition in the Jordan‐form representation.
Lemma 3.2.
@@
$\operatorname{rank}\left[\widehat{b}_{r_{i1}}^{i1\top},\widehat{b}_{r_{i2}}^{i2\top},\cdots,\widehat{b}_{r_{i\alpha_{i}}}^{i\alpha_{i}\top}\right]=\alpha_{i},\forall i=1,2,\cdots,l,$ (13)
where $\alpha_{i}$ is the geometric multiplicity of eigenvalue $\lambda_{i}$ of matrix $\overline{A}$. This means the last rows of $\widehat{{B}}_{i1},\widehat{{B}}_{i2},\cdots,\widehat{{B}}_{i\alpha_{i}}$ are linearly independent.
However, if there is a geometric multiplicity $\alpha_{i}$ exceeding the number of columns of $\widehat{B}$, i.e., $\exists\alpha_{i}>m$, it is impossible for the system to be fully controllable. In this case, we give a construction of $\widehat{B}$ which maximizes the controllable subspace.
@@
For $\widehat{B}$, we assume that the last rows of $\widehat{{B}}_{i1},\widehat{{B}}_{i2},\cdots,\widehat{{B}}_{i\alpha_{i}}$ satisfy
$\left[\!\begin{array}[]{c}\widehat{b}_{r_{i1}}^{i1}\\ \widehat{b}_{r_{i2}}^{i2}\\ \vdots\\ \widehat{b}_{r_{i\alpha_{i}}}^{i\alpha_{i}}\end{array}\!\right]\!\!=\!\!\left\{\begin{array}[]{cc}\left[I_{\alpha_{i}}\ \textbf{0}_{\alpha_{i}\times\left(m-\alpha_{i}\right)}\right],\,\text{for}\,\{i\mid m\geq\alpha_{i}\},\\ \\ \left[\delta_{\alpha_{i}}^{j_{1}},\delta_{\alpha_{i}}^{j_{2}},\cdots,\delta_{\alpha_{i}}^{j_{m}}\right],\,\text{for}\,\{i\mid m<\alpha_{i}\},\end{array}\right.$ (14)
where  $\delta_{\alpha_{i}}^{j_{p}}=\operatorname{Col}_{j_{p}}\left(I_{\alpha_{i}}\right)$, $\{j_{1},j_{2},\cdots,j_{m}\}\subset\{1,2,\cdots,\alpha_{i}\}$ are the subscripts of $m$-th largest Jordan blocks for eigenvalue $\lambda_{i}$ and other rows of $\widehat{{B}}_{i1},\widehat{{B}}_{i2},\cdots,\widehat{{B}}_{i\alpha_{i}}$ are zero rows. Then the controllable subspace is maximized.
Proof 3.4.
According to Lemma 3.2, we can easily derive that $\widehat{B}\in\mathbb{R}^{n\times m}$ maximizes the dimension of the controllable subspace if and only if the last rows of $\widehat{{B}}_{i1},\widehat{{B}}_{i2},\cdots,\widehat{{B}}_{i\alpha_{i}}$ satisfy
$\left\{\begin{array}[]{cc}\operatorname{rank}\left[\widehat{b}_{r_{i1}}^{i1\top},\widehat{b}_{r_{i2}}^{i2\top},\cdots,\widehat{b}_{r_{i\alpha_{i}}}^{i\alpha_{i}\top}\right]=\alpha_{i},\,\text{for}\,\{i\mid m\geq\alpha_{i}\},\\ \\ \operatorname{rank}\left[\widehat{b}_{r_{ij_{1}}}^{ij_{1}\top},\widehat{b}_{r_{ij_{2}}}^{ij_{2}\top},\cdots,\widehat{b}_{r_{ij_{m}}}^{ij_{m}\top}\right]=m,\,\text{for}\,\{i\mid m<\alpha_{i}\},\end{array}\right.$ (15)
where $\{j_{1},j_{2},\cdots,j_{m}\}\subset\{1,2,\cdots,\alpha_{i}\}$ are the subscripts of $m$-th largest Jordan blocks for eigenvalue $\lambda_{i}$. For $m\geq\alpha_{i}$, $\operatorname{rank}\left[I_{\alpha_{i}}\ \textbf{0}_{\alpha_{i}\times\left(m-\alpha_{i}\right)}\right]=\alpha_{i}$; for $m<\alpha_{i}$, $\operatorname{rank}\left[\delta_{\alpha_{i}}^{j_{1}},\delta_{\alpha_{i}}^{j_{2}},\cdots,\delta_{\alpha_{i}}^{j_{m}}\right]=m$. Thus the construction of $\widehat{B}$ in equation (14) maximizes the dimension of controllable subspace. This completes the proof.
With $\widehat{B}$ which satisfies Proposition 3.3, the best response for the defender is
@@
where $T$ is the transformation matrix making $(N+B_{a}^{\prime}U_{a})^{\top}$ become Jordan normal form. The value function with ${H}^{*}\in\mathbb{R}^{m\times n}$ is
$&\min_{{H}\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H)=\operatorname{dim\,Ker}\Omega(U_{a},H^{*})\!=\!n\!-\!\underset{{q}\in\mathcal{I}}{\min}(\sum_{p=1}^{m}r_{qj_{p}}),\,\text{for }\mathcal{I}\neq\emptyset,\\ &\min_{{H}\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H)=\operatorname{dim\,Ker}\Omega(U_{a},H^{*})=0,\,\text{for }\mathcal{I}=\emptyset,\end{aligned}\right.$ (16)
where =$\mathcal{I}=\{i\mid\alpha_{i}>m\}$ includes subscripts for the eigenvalues of $(N+B_{a}^{\prime}U_{a})$ whose geometric multiplicity is larger than $m$, $\{r_{qj_{1}},r_{qj_{2}},\cdots,r_{qj_{m}}\}$ are the dimensions of $m$-th largest Jordan blocks for eigenvalue $\lambda_{q}$. The calculation steps of $H^{*}$ based on Proposition 3.3 is summarized in Algorithm 2.
Algorithm 2 Defender: minimization of the unobservable subspace
  Input system matrices $N$, $B_{a}^{\prime}$; the attacker action $U_{a}$
  Set $\overline{A}=(N+B_{a}^{\prime}U_{a})^{\top}.$ Compute the Jordan normal form of $\overline{A}:J=T^{-1}\overline{A}T$, whose distinct eigenvalues are $\lambda_{1},\dots,\lambda_{l}$.
  for $i=1$ to $l$ do
     $\alpha_{i}\leftarrow$ geometric multiplicity of $\lambda_{i}$;
     $\{r_{i1},r_{i2},\cdots,r_{i\alpha_{i}}\}\leftarrow$ dimensions of Jordan blocks of $\lambda_{i}$;
     $\{j_{1},j_{2},\cdots,j_{m}\}\leftarrow$ the subscripts of $m$-th largest Jordan blocks of $\lambda_{i}$.
  end forCompute $\widehat{B}$ according to Corollary 3.3. $H^{*}=\widehat{B}^{\top}T^{\top}$.return $H^{*}$
@@ 4 Equilibrium analysis
Based on the above best response sets, we next give the equilibrium of the game considering one-step and two-step optimality respectively in subsections 4.1 and 4.2. Then in subsections 4.3 and 4.4, game outcomes and equilibrium characterization are refined, which are suitable for both one-step or two-step optimality. Finally, three key insights of the observability-adversarial game are summarized in subsection 4.5.
@@ 4.1 Game based on one-step optimality
In the one‐step optimality formulation of the repeated game, each player maximizes only the immediate utility at each stage. We denote the resulting best response sets for the attacker and the defender by $BR1^{a}$ and $BR1^{d}$, i.e.,
$BR1^{a}(H)=\arg\max_{{U_{a}}\in\mathbb{R}^{k\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H),\,BR1^{d}(U_{a})=\arg\min_{{H}\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H).$ (17)
The solutions of these best response sets have been discussed in subsections 3.1 and 3.2, i.e.,
$BR1^{a}(H)=\mathcal{F}(\mathcal{V}^{*}(H)),$ (18)
where $\mathcal{F}(\mathcal{V}^{*})$ is the friend matrix set of $\mathcal{V}^{*}$, $\mathcal{V}^{*}$ is the maximal controlled invariant subspace in $\operatorname{Ker}H$;
@@
where $T$ is the transformation matrix making $(N+B_{a}^{\prime}U_{a})^{\top}$ become Jordan normal form. Figure 2 shows the sequence of actions.
Figure 2  Sequence of actions considering one-step optimality.
Both best response sets are not single-valued maps and Algorithm 1 (or Algorithm 2) only chooses a special $U_{a}$ (or $H$) in $BR1^{a}$ (or $BR1^{d}$). If a player chooses a different action in the best response set, the game result will be different.
Define the Nash equilibrium based on one-step optimality as follows.
Definition 4.1.
The strategy profile $(U_{a}^{*},H^{*})$ is said to be the Nash equilibrium (NE) of the one-step optimal game, if $U_{a}^{*}\in BR1^{a}(H^{*}),H^{*}\in BR1^{d}(U_{a}^{*})$.
@@
Theorem 4.2 (One-step optimality NE criterion).
The strategy profile $(U_{a}^{*},H^{*})$ is a Nash equilibrium of the one-step optimal game if and only if
$\min_{{H}\in\mathbb{R}^{m\times n}}\!\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},\!H)=\operatorname{dim}\mathcal{V}^{*}(H^{*}).$ (20)
Proof 4.3.
(Sufficiency) We prove $U_{a}^{*}\in BR1^{a}(H^{*}),H^{*}\in BR1^{d}(U_{a}^{*})$ by establishing its contrapositive. If $U_{a}^{*}\notin BR1^{a}(H^{*})$, there is $\operatorname{dim\,Ker}\Omega(U_{a}^{*},H^{*})<\operatorname{dim}\mathcal{V}^{*}(H^{*})$. Because there is $\min_{{H}\in\mathbb{R}^{m\times n}}\!\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},\!H)\leq\operatorname{dim\,Ker}\Omega(U_{a}^{*},H^{*})$, this contradicts (20). If $H^{*}\notin BR1^{d}(U_{a}^{*})$, there is $\min_{{H}\in\mathbb{R}^{m\times n}}\!\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},\!H)<\operatorname{dim\,Ker}\Omega(U_{a}^{*},H^{*})$. Because there is $\operatorname{dim\,Ker}\Omega(U_{a}^{*},H^{*})$ $\leq\operatorname{dim}\mathcal{V}^{*}(H^{*})$, this contradicts (20). Thus there must be $U_{a}^{*}\in BR1^{a}(H^{*}),H^{*}\in BR1^{d}(U_{a}^{*})$ and $(U_{a}^{*},H^{*})$ is a Nash equilibrium.
(Necessity) Because $U_{a}^{*}\in BR1^{a}(H^{*})$, there is
$\operatorname{dim\,Ker}\Omega(U_{a}^{*},H^{*})\,=\operatorname{dim}\mathcal{V}^{*}(H^{*}).$ (21)
Since $H^{*}\in BR1^{d}(U_{a}^{*})$, we have
$\min_{{H}\in\mathbb{R}^{m\times n}}\!\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},\!H)=\operatorname{dim\,Ker}\Omega(U_{a}^{*},H^{*}).$ (22)
Combining (21) and (22), we can get (20), which completes the proof.
Theorem 4.2 is effective in determining when the Nash equilibrium is reached in dynamic game process, as it only requires verifying whether the value function $\operatorname{dim\,Ker}\Omega\,(U_{a}^{*},\!H)$ after the defender chooses $H^{*}$ is equal to $\operatorname{dim}\mathcal{V}^{*}(H^{*})$.
4.2 Game based on two-step optimality
Since the one-step best response sets, are multi-valued, different preferences for choices in $BR1^{a}$ and $BR1^{d}$ can steer the play toward different result. For any $H\in BR1^{d}$, the induced value $\dim\mathcal{V}^{*}$ varies, thereby tightening or relaxing the upper bound of $\max_{U_{a}}\dim\operatorname{Ker}\Omega(U_{a},H)$. Conversely, every $U_{a}\in BR1^{a}$ alters the maximum geometric multiplicity among all eigenvalues of $N+B_{a}^{\top}U_{a}$, which fixes the lower bound of $\min_{H}\dim\operatorname{Ker}\Omega(U_{a},H)$.
To capture these continuation effects, we impose a two-step optimality criterion: within $BR1^{a}$ or $BR1^{d}$, each player selects the action that maximizes the loss of the opponent in the subsequent stage. The attacker’s two-step best response set is
$BR2^{a}(H)=\arg\max_{U_{a}\in BR1^{a}(H)}\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=\arg\max_{U_{a}\in BR1^{a}(H)}[\text{MGM}(N+B_{a}^{\prime}U_{a})].$ (23)
[…]
$BR2^{d}(U_{a})=arg\min_{H\in BR1^{d}(U_{a})}\max_{U_{a}\in\mathbb{R}^{k\times n}}\operatorname{dim\,Ker}\Omega\,(U_{a},H)=arg\min_{H\in BR1^{d}(U_{a})}\operatorname{dim\,}\mathcal{V}^{*}(H),$ (24)
which can be derived from (10).
For comparison, define $BR2X^{a}\triangleq\arg\max_{U_{a}\in\mathbb{R}^{k\times n}}[\text{MGM}(N+B_{a}^{\prime}U_{a})]$, $BR2X^{d}\triangleq\arg\,$ $\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H)$. Neither $BR2X^{a}$ nor $BR2X^{d}$ is a valid two‐step best response, since each ignores the requirement to optimize the current period’s value function. Instead, they correspond exactly to the Stackelberg solutions under two different leadership orders: when the defender leads, it commits to $BR2X^{d}$ and the attacker responds with $BR1^{a}$; when the attacker leads, it chooses $BR2X^{a}$ and the defender replies with $BR1^{d}$.
Figure 3 shows the sequence of actions following two-step optimality.
Figure 3  Sequence of actions following two-step optimality.
Then we define the Nash equilibrium when two players consider two-step optimality.
Definition 4.4.
The strategy profile $(U_{a}^{*},H^{*})$ is said to be a Nash equilibrium of the two-step optimal game if $U_{a}^{*}\in BR2^{a}(H^{*}),H^{*}\in BR2^{d}(U_{a}^{*})$.
In order to get the condition for Nash equilibrium under two-step optimality, we give the following lemma.
Lemma 4.5.
For the same $N\in\mathbb{R}^{n\times n},B_{a}^{\prime}\in\mathbb{R}^{n\times k}$, $\forall U_{a}\in\mathbb{R}^{k\times n}$, there is
$\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,}\mathcal{V}^{*}(H)\geq\min_{H\in\mathbb{R}^{m\times n}}\operatorname{dim\,Ker}\Omega(U_{a},H).$ (25)